OpenStudy (anonymous):

Use the Factor Theorem to determine if (x - 4) is a factor of x3 + x2 - 16x - 16 A. Factor B. Not a factor PLEASEE HELPPP IM SO SLOW

ganeshie8 (ganeshie8):

divide (x-4) and see wat u get as remainder

ganeshie8 (ganeshie8):

if you get 0 remainder, then (x-4) is a factor.

OpenStudy (anonymous):

I still do not understand what am i dividing (x-4) too?????????

OpenStudy (anonymous):

ohhhhhhhhhhhh nvm i get it

ganeshie8 (ganeshie8):

ohk... good :) so wat did u get as remainder ?

OpenStudy (anonymous):

0 was the remainder

ganeshie8 (ganeshie8):

Yes ! gw !!

OpenStudy (anonymous):

Given the hyperbola (x2/4) – (y2/9) = 1, find the equations for its asymptotes. A. x = (1/3)y, x = (–1/3)y B. y = (1/3)x, y = (–1/3)x C. y = (3/2)x, y = (–3/2)x D. x = (3/2)y, x = (–3/2)y

OpenStudy (anonymous):

HELP LOL

ganeshie8 (ganeshie8):

not good wid hyperbolas.... :( post in a new q.... somebody wil help :)

OpenStudy (anonymous):

I DID BUT NOBODY IS LOL

OpenStudy (anonymous):

sorry caps

ganeshie8 (ganeshie8):

its okay..... wait some time... someone will :)

OpenStudy (anonymous):

lol ughhh

OpenStudy (anonymous):

What is the degree of the monomial 32ab2c? A. 1 B. 2 C. 3 D. 4